# SGS-SLAM: Semantic Gaussian Splatting for Neural Dense SLAM

### 25 Feb 2024 v2

@@ 4.1.1 Datasets
We evaluate our method on both synthetic and real-world datasets. To compare with other neural implicit SLAM methods, we evaluate on 8 synthetic scenes from Replica dataset Straub et al. (2019) and real-world scenes from ScanNet Dai et al. (2017a) and ScanNet++ Yeshwanth et al. (2023) datasets. The ground-truth camera pose and semantic map of Replica are offered from simulation, and ground-truth camera pose of ScanNet is generated by BundleFusion Dai et al. (2017b). The ground-truth semantic label is provided by the dataset.
@@ 4.1.2 Metrics
We use PSNR, Depth-L1 (on 2D depth map), SSIM, and LPIPS to evaluate the reconstruction quality. For the evaluation of camera pose, we adopt the average absolute trajectory error (ATE RMSE). For semantic segmentation, we calculate mIoU score.
@@ 4.1.3 Baselines
We compare the tracking and mapping with state-of-the-art methods iMap Sucar et al. (2021), Vox-Fusion Yang et al. (2022), NICE-SLAM Zhu et al. (2022), Co-SLAM Wang et al. (2023), ESLAM Johari et al. (2023), and SplaTAM Keetha et al. (2023). For semantic segmentation accuracy, we compare with NIDS-SLAM Haghighi et al. (2023), DNS-SLAM Li et al. (2023a), and SNI-SLAM Zhu et al. (2023).
@@ 4.2 Evaluation of Mapping and Localization
We show quantitative measures of reconstruction quality using the Replica dataset Straub et al. (2019) in table 1. Our method demonstrates state-of-the-art performance. When compared to other baseline methods, our approach attains notably superior outcomes, outperforming them by a margin of 10dB in PSNR.
@@
In fig. 2, we present the reconstruction results of three chosen scenes, where regions of interest are accentuated with boxes in various colors. Our method exhibits high-fidelity reconstruction outcomes. Specifically, for small, intricately textured objects like a clock, socket, books on a tea table, and a lamp, our approach shows remarkable accuracy over NeRF-based methods. This is because Gaussians are capable of representing objects with complex textures and surfaces. Furthermore, NeRF-based methods often struggle with the over-smoothing issue, resulting in blurred edges on objects. In contrast, by utilizing an explicit Gaussian representation, SGS-SLAM precisely captures objects with clear edges, irrespective of their sizes. Compared with SplaTAM, which is also a Gaussian-based model, our approach utilizes semantic information for discerning object categories, recognizing visual appearance to determine texture, and applies geometric constraints to preserve accurate shapes. This combination enables our method to achieve thorough modeling of both objects and their surrounding environment. The combination of these constraints allows SGS-SLAM to capture fine-grained details of objects, offering high-fidelity and accurate reconstruction.
@@ 4.3 Evaluation of Semantic Segmentation
table 3 shows a quantitative evaluation of our method in comparison to other neural semantic SLAM approaches. It’s worth noting that we only show four scenes because previous NeRF-based semantic models only reported results on these scenes. In comparison to these previous methods, SGS-SLAM demonstrates state-of-the-art performance, outperforming the initial baseline by more than 10%. Substantial enhancement highlights the crucial advantage of explicit Gaussian representation over NeRF-based aprroaches. Gaussians can precisely isolate object boundaries, resulting in highly accurate 3D scene segmentation. In contrast, NeRF-based methods often struggle to recognize individual objects and typically require complex muti-level model designs and extensive feature fusion. Our approach offers an unparalleled ability to identify 3D objects in decomposed representations, which can serve as 3D priors for tracking and mapping in future time steps, and is well-suited for further downstream tasks.
@@ 4.4 Evaluation of Key Frame Optimization
Figure 3: The selected novel view evaluation of scene0000 from the ScanNet dataset Dai et al. (2017a). The rendered views display the reconstructed bike and guitar captured within the trajectory. Our method outperforms baselines by a large margin primarily due to the integration of key-frame optimization and semantic constraints. Note that the ground-truth for novel views is derived from the pre-established mesh provided by the ScanNet dataset.
In real-world datasets, tracking errors tend to accumulate along a trajectory, making pose estimations at later timestamps less reliable. Such inaccuracies can compromise the quality of map reconstructions, negatively impacting the accuracy of a previously well-established scene. A case in point is scene0000 from the ScanNet dataset Dai et al. (2017a), where objects such as bike and guitar are revisited at early and late stages in the trajectory. Key frames from later in the sequence, influenced by inaccurate camera poses, can disrupt the previously accurate reconstructions of these objects. fig. 3 illustrates the novel-view evaluation for scene0000. In comparison to SplaTAM, depicted in the upper row, our method delivers more accurate reconstruction outcomes. The bike, garbage bin, and guitar are accurately rendered, meanwhile details are preserved. Our method facilitates the selection of key frames base on geometric and semantic constraints, incorporating an uncertainty weighting during the optimization of selected key frames. This strategy demonstrate its effectiveness in map optimization from different views meanwhile preventing the unreliable key frame with high uncertainty to significantly altering the earlier accurately reconstructed map.
@@ 4.5 Scene Manipulation
The obtained semantic mask within the 3D scene has a range of applications for subsequent tasks. As an illustrative example, we demonstrate an straightforward but efficient Gaussian editing method defined by eq. 10, which is crucial for enabling scene manipulation for robotics or mixed reality applications.
Figure 4: The case study on scene manipulation in room0 of the Replica dataset Straub et al. (2019). We show the capabilities for object removal and transformation by specifying semantic labels. SGS-SLAM allows manipulation of either individual objects or a group of items, as illustrated by actions that include the removal of a table, as well as moving and rotating the table together with all objects on it.
Utilizing the decoupled scene representation, in contrast to NeRF-based approaches that demand fine-tuning of the entire network, SGS-SLAM has the ability to choose specific objects within the scene by referencing their semantic mask while keeping the remainder of the well-trained, irrelevant environment fixed. As shown in fig. 4, for object removal, we can directly erase the Gaussians associated with the editing target, such has removing the table while preserving all the items on it. In addition, we can group objects by selecting their semantic masks and apply translation and rotation, such as moving and rotating both the table and the above objects to a different place. This editing capability requires no training or fine-tuning, making it readily available for downstream applications.
@@ 4.6 Ablation Study
We perform the ablation of GS-SLAM on the scene0000 of the ScanNet dataset Dai et al. (2017a) to evaluate the effectiveness of multi-channel feature supervision, and the key-frame optimization strategies.
@@ 4.6.1 Effect of Multi-channel Optimization
table 4 shows the ablation study on multi-channel parameter optimization employed in SGS-SLAM. The results reveal that our optimization strategy can significantly improve the localization and mapping performance. Specifically, the system without appearance color, although performs poorly, still maintains its capability for tracking and depth estimation. This can be attributed to the gradient flow resulting from semantic supervision. Despite the Gaussian rasterizer used in our system not supporting backward flow through depth rendering, this underscores the value of semantic supervision within our framework. The absence of depth data leads to the poorest depth estimation, highlighting the importance of geometric supervision. Furthermore, the absence of an input semantic map disables 3D semantic segmentation and remarkably diminishes the performance of tracking and mapping. Additionally, the silhouette threshold, essential for assessing scene visibility, is crucial for the system stability. Without this threshold, the system shows a significant decline in the effectiveness of tracking and mapping.
@@
w/o color image 7.44 24.59 68.19
w/o depth map 47.66 40.47 15.14 54.52
w/o semantic map 9.15 13.81 17.52
w/o silhouette threshold 29.12 357.48 12.06 28.07
w/ multi-channel inputs 6.66 11.26 19.47 70.27
4.6.2 Effect of Key-frame Optimization
table 5 presents the results of key-frame selection ablation. Our two-level key-frame selection strategy reveals that omitting either geometric or semantic constraints results in a significant drop in both tracking and mapping performance. Additionally, without incorporating uncertainty weighting, the system demonstrates decrease in performance compared to its full implementation.
Table 5: Ablation study of key-frame optimization on the scene0000 of the ScanNet dataset Dai et al. (2017a). The comparison involves settings where geometric, semantic, and uncertainty constraints are removed.
@@ 4.6.2 Effect of Key-frame Optimization
w/o geometric threshold 6.66 15.55 19.21 68.93
w/o semantic threshold 8.44 12.89 17.84 69.85
w/o uncertainty weighting 6.87 11.43 18.72 70.12
w/ key-frame selection 6.18 11.26 19.47 70.27
@@ 5 Conclusion and Limitations
We presented SGS-SLAM, the first semantic dense visual SLAM system based on the 3D Gaussian representation. We propose to leverage multi-channel parameter optimization where appearance, geometric, and semantic constraints are combined to enforce high-accurate 3D semantic segmentation, and high-fidelity dense map reconstruction meanwhile effectively produce a robust camera pose estimation. SGS-SLAM takes the advantages of optimal key-frame optimization, resulting on reliable reconstruction quality. Extensive experiments show that our method provide state-of-the-art tracking and mapping results, meanwhile maintain rapid rendering speeds. Furthermore, the high-quality reconstruction of scenes and precise 3D semantic labeling generated by our system establish a strong foundation for the downstream tasks such as scene editing, offers solid prior for the robotics or AR/VR applications.
@@ 5.1 Limitations
SGS-SLAM replies on depth and 2D semantic signal inputs for tracking and mapping. In scenarios where this information is scarce or difficult to access, the system’s effectiveness will be compromised. Additionally, our method incurs large memory consumption when deployed in large scenes. Addressing these limitations will be an objective for future research.
@@ Appendix A Experiment Settings
In this section, we outline the experimental setup and hyperparameters applied in our studies.
A.1 SGS-SLAM Settings
By default, both mapping and tracking operations are conducted for each frame. During the tracking phase, we set the silhouette visibility threshold, $T_{\rm sil}$, to 0.99. The multi-channel optimization involves three parameters: $\lambda_{D}=1.0$ for depth, $\lambda_{C}=0.5$ for colors, and $\lambda_{S}=0.05$ for semantic loss, with the semantic loss weight being comparatively low due to the typical noisiness of real-world semantic labels. Throughout the tracking, the multi-channel Gaussian parameters remain constant, adjusting only the camera parameters with a learning rate of 2e-3 for transition. Key-frames are initially chosen at intervals of every 5 frames, then refined based on geometric and semantic criteria. The geometric overlap threshold, $\eta$, is defined at 0.05, and the semantic mean Intersection over Union (mIoU) threshold, $T_{\rm sem}$, at 0.7. The maximum number of key-frames per frame is limited to 25, considering the computation speed. The uncertainty decay coefficient, $\tau$ scales with the length of the input frame series. In mapping process, the silhouette threshold $T_{\rm sil}$ is adjusted to 0.5. The weights of photometric loss are set to $\lambda_{D}=1.0$, $\lambda_{C}=0.5$, and $\lambda_{S}=0.1$. Here, camera parameters are fixed, and Gaussian parameters are optimized, with specific learning rates for 3D position at 1e-4, color 2.5e-3, Gaussian rotation at 1e-3, logit opacity at 0.05, and logg scale at 1e-3. Performance metrics of tracking and mapping are assessed every 5 frames, with mIoU scores evaluated at the same frequency.
@@ B.2 Semantic Segmentation
Figure 5: Qualitative comparison of our method and DNS-SLAM Li et al. [2023a] for semantic segmentation from the Replica dataset Straub et al. [2019]. The visualization outcomes of DNS-SLAM Li et al. [2023a] are obtained from its paper. The frames of training view are chosen based on the results presented in DNS-SLAM. Compared to NeRF-based models, our approach delivers segmentation results with higher accuracy.
@@ B.3 Novel View Rendering
Figure 7: The visualization of novel view rendering between the baseline and our method using the ScanNet dataset Dai et al. [2017a]. The ground-truth novel views are captured from meshes. SGS-SLAM exhibits rendering of high fidelity, outperforms the NeRF-based ESLAM Johari et al. [2023]. In contrast to the ground-truth mesh, our method demonstrates robust mapping in areas where the ground-truth mesh presents holes.
Figure 8: The visualization of 3D semantic segmentation results of SGS-SLAM, as applied to the novel views selected in fig. 7. Note that the rendering results exhibit minor variations in scene objects due to the use of an modified semantic dataset from ScantNet. For our method, the training data is processed from the filtered semantic labels using the nyu40-class, where certain objects are not distinctly labeled and are assigned as background (depicted in black). Furthermore, we introduce extra labels, like guitar, bag, and basket, to enhance the quality of scene reconstruction.

### 02 Mar 2024 v3

@@ 4.1.1 Datasets
We evaluate our method on both synthetic and real-world datasets. To compare with other neural implicit SLAM methods, we evaluate synthetic scenes from Replica dataset Straub et al. (2019) and real-world scenes from ScanNet Dai et al. (2017a) and ScanNet++ Yeshwanth et al. (2023) datasets. The ground-truth camera pose and semantic map of Replica are offered from simulation, and the ground-truth camera pose of ScanNet is generated by BundleFusion Dai et al. (2017b). The ground-truth 2D semantic label is provided by the dataset.
@@ 4.1.2 Metrics
We use PSNR, Depth-L1 (on 2D depth maps), SSIM, and LPIPS to evaluate the reconstruction quality. For the evaluation of camera pose, we adopt the average absolute trajectory error (ATE RMSE). For semantic segmentation, we calculate mIoU score.
@@ 4.1.3 Baselines
We compare the tracking and mapping with state-of-the-art methods iMap Sucar et al. (2021), NICE-SLAM Zhu et al. (2022), Co-SLAM Wang et al. (2023), ESLAM Johari et al. (2023), and SplaTAM Keetha et al. (2023). For semantic segmentation accuracy, we compare with NIDS-SLAM Haghighi et al. (2023), DNS-SLAM Li et al. (2023a), and SNI-SLAM Zhu et al. (2023).
@@ 4.2 Evaluation of Mapping and Localization
We show quantitative measures of reconstruction quality using the Replica dataset Straub et al. (2019) in table 1. Our method demonstrates state-of-the-art performance. Compared to other baseline methods, our approach attains notably superior outcomes, outperforming them by a margin of 10dB in PSNR.
@@
In fig. 2, we present the reconstruction results of three chosen scenes, where regions of interest are accentuated with boxes in various colors. Our method exhibits high-fidelity reconstruction outcomes. Specifically, for small, intricately textured objects like a clock, socket, books on a tea table, and a lamp, our approach shows remarkable accuracy over NeRF-based methods. This is because Gaussians are capable of representing objects with complex textures and surfaces. Furthermore, NeRF-based methods often struggle with the over-smoothing issue, resulting in blurred edges on objects. In contrast, by utilizing an explicit Gaussian representation, SGS-SLAM precisely captures objects with clear edges, irrespective of their sizes. Compared with SplaTAM, which is also a Gaussian-based model, our approach utilizes semantic information for discerning object categories, recognizing visual appearance to determine texture, and applying geometric constraints to preserve accurate shapes. This combination enables our method to achieve thorough modeling of both objects and their surrounding environment. The combination of these constraints allows SGS-SLAM to capture fine-grained details of objects, offering high-fidelity and accurate reconstruction.
@@ 4.3 Evaluation of Semantic Segmentation
table 3 shows a quantitative evaluation of our method in comparison to other neural semantic SLAM approaches. It’s worth noting that we only show four scenes because previous NeRF-based semantic models only reported results on these scenes. In comparison to these previous methods, SGS-SLAM demonstrates state-of-the-art performance, outperforming the initial baseline by more than 10%. Substantial enhancement highlights the crucial advantage of explicit Gaussian representation over NeRF-based approaches. Gaussians can precisely isolate object boundaries, resulting in highly accurate 3D scene segmentation. In contrast, NeRF-based methods often struggle to recognize individual objects and typically require complex muti-level model designs and extensive feature fusion. Our approach offers an unparalleled ability to identify 3D objects in decomposed representations, which can serve as 3D priors for tracking and mapping in future time steps, and is well-suited for further downstream tasks.
@@ 4.4 Evaluation of Key Frame Optimization
Figure 3: The selected novel view evaluation of scene0000 from the ScanNet dataset Dai et al. (2017a). The rendered views display the reconstructed bike and guitar captured within the trajectory. Our method outperforms baselines by a large margin primarily due to the integration of keyframe optimization and semantic constraints. Note that the ground-truth for novel views is derived from the pre-established mesh provided by the ScanNet dataset.
In real-world datasets, tracking errors tend to accumulate along a trajectory, making pose estimations at later timestamps less reliable. Such inaccuracies can compromise the quality of map reconstructions, negatively impacting the accuracy of a previously well-established scene. A case in point is scene0000 from the ScanNet dataset Dai et al. (2017a), where objects such as bike and guitar are revisited at early and late stages in the trajectory. Keyframes from later in the sequence, influenced by inaccurate camera poses, can disrupt the previously accurate reconstructions of these objects. fig. 3 illustrates the novel-view evaluation for scene0000. In comparison to SplaTAM, depicted in the upper row, our method delivers more accurate reconstruction outcomes. The bike, garbage bin, and guitar are accurately rendered, meanwhile details are preserved. Our method facilitates the selection of keyframes based on geometric and semantic constraints, incorporating uncertainty weighting during the optimization of selected keyframes. This strategy demonstrates its effectiveness in map optimization from different views meanwhile preventing the unreliable key frame with high uncertainty to significantly altering the earlier accurately reconstructed map.
@@ 4.5 Scene Manipulation
The obtained semantic mask within the 3D scene has a range of applications for subsequent tasks. As an illustrative example, we demonstrate a straightforward but efficient Gaussian editing method defined by eq. 10, which is crucial for enabling scene manipulation for robotics or mixed reality applications.
Figure 4: The case study on scene manipulation in room0 of the Replica dataset Straub et al. (2019). We show the capabilities for object removal and transformation by specifying semantic labels. SGS-SLAM allows manipulation of either individual objects or a group of items, as illustrated by actions that include the removal of a jar and flowers, as well as moving and rotating them.
Utilizing the decoupled scene representation, in contrast to NeRF-based approaches that demand fine-tuning of the entire network, SGS-SLAM can edit specific objects within the scene while keeping the remainder of the well-trained, irrelevant environment fixed. As shown in fig. 4, we can directly manipulate the Gaussians associated with the editing target, such as erasing, moving, and rotating the jar and flowers on the table. In addition, we can group objects by selecting their semantic masks and applying a transition, such as rotating both the table and the above objects as shown in the supplementary material. This editing capability requires no training or fine-tuning, making it readily available for downstream applications.
@@ 4.6 Ablation Study
We perform the ablation of GS-SLAM on the scene0000 of the ScanNet dataset Dai et al. (2017a) to evaluate the effectiveness of multi-channel feature supervision, and the keyframe optimization strategies.
@@ 4.6.1 Effect of Multi-channel Optimization
table 4 shows the ablation study on multi-channel parameter optimization. The results reveal that our optimization strategy can significantly improve the localization and mapping performance. Specifically, the system without appearance color cannot provide rendered views, whereas camera pose and depth can still be estimated by leveraging depth and semantic input. The absence of depth data leads to the poorest depth estimation, highlighting the importance of geometric supervision. Furthermore, the absence of an input semantic map disables 3D semantic segmentation and remarkably diminishes the performance of tracking and mapping. Additionally, the silhouette threshold, essential for assessing scene visibility, is crucial for the system stability. Without this threshold, the system shows a significant decline in the effectiveness of tracking and mapping.
@@
without color image 7.44 24.59 68.19
without depth map 47.66 40.47 15.14 54.52
without semantic map 9.15 13.81 17.52
without silhouette threshold 29.12 357.48 12.06 28.07
with multi-channel inputs 6.66 11.26 19.47 70.27
4.6.2 Effect of Keyframe Optimization
table 5 presents the results of keyframe selection ablation. Our two-level keyframe selection strategy reveals that omitting either geometric or semantic constraints results in a significant drop in both tracking and mapping performance. Additionally, without incorporating uncertainty weighting, the system demonstrates a decrease in performance compared to its full implementation.
Table 5: Ablation study of keyframe optimization on the scene0000 of the ScanNet dataset Dai et al. (2017a). The comparison involves settings where geometric, semantic, and uncertainty constraints are removed.
@@ 4.6.2 Effect of Keyframe Optimization
without geometric threshold 6.66 15.55 19.21 68.93
without semantic threshold 8.44 12.89 17.84 69.85
without uncertainty weighting 6.87 11.43 18.72 70.12
with keyframe selection 6.18 11.26 19.47 70.27
@@ 5 Conclusion and Limitations
We presented SGS-SLAM, the first semantic dense visual SLAM system based on the 3D Gaussian representation. We propose to leverage multi-channel parameter optimization where appearance, geometric, and semantic constraints are combined to enforce high-accurate 3D semantic segmentation, and high-fidelity dense map reconstruction meanwhile effectively producing a robust camera pose estimation. SGS-SLAM takes advantage of optimal keyframe optimization, resulting in reliable reconstruction quality. Extensive experiments show that our method provides state-of-the-art tracking and mapping results, meanwhile maintaining rapid rendering speeds. Furthermore, the high-quality reconstruction of scenes and precise 3D semantic labeling generated by our system establish a strong foundation for downstream tasks such as scene editing, offering solid prior for robotics or mixed reality applications.
@@ 5.1 Limitations
SGS-SLAM replies on depth and 2D semantic signal inputs for tracking and mapping. In scenarios where this information is scarce or difficult to access, the system’s effectiveness will be compromised. Additionally, our method incurs large memory consumption when deployed to large scenes. Addressing these limitations will be an objective for future research.
@@ Appendix A Experiment Settings
In this section, we outline the experimental setup and hyperparameters applied in our studies. The experiments were conducted on NVIDIA A100-40GB GPUs.
A.0.1 SGS-SLAM
By default, both mapping and tracking operations are conducted for each frame. During the tracking phase, we set the silhouette visibility threshold, $T_{\rm sil}$, to 0.99. The multi-channel optimization involves three parameters: $\lambda_{D}=1.0$ for depth, $\lambda_{C}=0.5$ for colors, and $\lambda_{S}=0.05$ for semantic loss, with the semantic loss weight being comparatively low due to the typical noisiness of real-world semantic labels. Throughout the tracking, the multi-channel Gaussian parameters remain constant, adjusting only the camera parameters with a learning rate of 2e-3 for transition. Key-frames are initially chosen at intervals of every 5 frames, then refined based on geometric and semantic criteria. The geometric overlap threshold, $\eta$, is defined at 0.05, and the semantic mean Intersection over Union (mIoU) threshold, $T_{\rm sem}$, at 0.7. The maximum number of keyframes per frame is limited to 25, considering the computation speed. The uncertainty decay coefficient, $\tau$ scales with the length of the input frame series. In mapping process, the silhouette threshold $T_{\rm sil}$ is adjusted to 0.5. The weights of photometric loss are set to $\lambda_{D}=1.0$, $\lambda_{C}=0.5$, and $\lambda_{S}=0.1$. Here, camera parameters are fixed, and Gaussian parameters are optimized, with specific learning rates for 3D position at 1e-4, color 2.5e-3, Gaussian rotation at 1e-3, logit opacity at 0.05, and log scale at 1e-3. Performance metrics of tracking and mapping are assessed every 5 frames, with mIoU scores evaluated at the same frequency.
@@ B.2 Semantic Segmentation
Figure 5: Qualitative comparison of our method and DNS-SLAM Li et al. [2023a] for semantic segmentation from the Replica dataset Straub et al. [2019]. The visualization outcomes of DNS-SLAM Li et al. [2023a] are obtained from its paper. The frames of the training view are chosen based on the results presented in DNS-SLAM. Compared to NeRF-based models, our approach delivers segmentation results with higher accuracy.
@@ B.3 Novel View Rendering
Figure 7: The visualization of novel view rendering between the baseline and our method using the ScanNet dataset Dai et al. [2017a]. The ground-truth novel views are captured from meshes. SGS-SLAM exhibits rendering of high fidelity and outperforms the NeRF-based ESLAM Johari et al. [2023]. In contrast to the ground-truth mesh, our method demonstrates robust mapping in areas where the ground-truth mesh presents holes.
Figure 8: The visualization of 3D semantic segmentation results of SGS-SLAM, as applied to the novel views selected in fig. 7. Note that the rendering results exhibit minor variations in scene objects due to the use of a modified semantic dataset from ScantNet. For our method, the training data is processed from the filtered semantic labels using the nyu40-class, where certain objects are not distinctly labeled and are assigned as background (depicted in black). Furthermore, we introduce extra labels, like guitar, bag, and basket, to enhance the quality of scene reconstruction.
@@
B.4 Scene Manipulation
In this section, we visualize scene manipulation results by grouping the Gaussians using the semantic mask. As shown in fig. 10, for object removal, we can directly erase the Gaussians associated with the editing target, such as removing the table while preserving all the items on it. In addition, we can group objects by selecting their semantic masks and applying translation and rotation, such as moving and rotating both the table and the above objects to a different place.
Figure 10: The visualization of scene manipulation by grouping Gaussians via semantic labels. SGS-SLAM allows manipulation of either individual objects or a group of items, as illustrated by actions that include the removal of a table, as well as moving and rotating the table together with all objects on it.